How much quantity of 15% of acid solution must be added to 8L of 30% solution of the same acid to obtain 25% solution for duty cleaning?
Answer by mananth(16946)

15% of acid solution ---let it be x liters
30% solution --------8 liters
mixture 25% solution (8+x) liters
..
0.15x+0.08*30=0.25(8+x)
0.15x+2.4 =2+0.25x
add -2.4
0.15x+2.4-2.4 = 2-2.4 +0.25x
0.15x = 0.4+0.25x
0.1x= 0.4
x= 4 liters of 15% solution